The 2018 balance sheet of Speith’s Golf Shop, Inc., showed long-term debt of $2.5 million, and the 2019 balance sheet showed long-term debt of $2.75 million. The 2019 income statement showed an interest expense of $100,000. The 2018 balance sheet showed $450,000 in the common stock account and $2.1 million in the additional paid-in surplus account. The 2019 balance sheet showed $490,000 and $2.4 million in the same two accounts, respectively. The company paid out $550,000 in cash dividends during 2019. Suppose you also know that the firm’s net capital spending for 2019 was $1,310,000, and that the firm reduced its net working capital investment by $57,000.

What was the firm’s 2019 operating cash flow, or OCF?
Answer #1

C D E HNM + Long term debt Common stock additional paid-in surplus account. 2018 2019 2500000 2750000 450000 490000 2100000 2

Ham to 000 Long term debt Common stock additional paid-in surplus account. 2018 =2.5*10^6 450000 =2.1*10^6 2019 =2.75*10^6 49